A little header

me

February 4, 2013

This is a sample to show how to incorporate graphs within a latex document the basic idea is to produce the graph in (Encapsulated) Postscript using matlab or mathematica, and the import it here. The formats .ps and .eps will work, but .eps is preferred. The .dvi file produced by L^{A}TEX does not actually include the figures, but xdvi will usually render them. To actually include them, convert to postscript using dvips.

For this file to work properly, the file lfig.eps must be in the same directory. You can download it from the class web page. This file was created by matlab, using the “Export… ” option on the “File” menu of the Figure window.

1 a section

This text was written before Figure 1, which we request appear right after it.

Figure 1: adjusts the figure width and height

Then we have more text, and request Figure 2 appear at the top of a page. filler

filler filler filler filler filler filler filler filler filler filler filler filler filler filler filler filler filler filler

Figure 2: adjusts the figure width, letting the height change automatically

filler filler filler

You can alse do things like rotate and box (see Figure 3).

Figure 3: rotates

More text, and a big Figure 4, which we ask to go on a page with only figures. me 3

Figure 4: lets the figure be its natural size